Pacita was 25 years old when her first son was born. Today, the sum of their ages is 105. How old are they now?

Answer by addingup(3677)

let x be the mother and y the son:
x = y+25 (mother was 25 when son was born)
x+y = 105 substitute for x:
y+25+y = 105
2y = 80
y = 40 the son is now 40, and the mother:
x = 40+25 = 65 is the age of the mother
---------------------------
check;
40+65 = 105 correct
:
John